# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 8 Results
DR3 is the first release providing orbital solutions for the RV time-series. As such, the present results must be considered preliminary even if the performances are already good.
@@ 8.1 Tallies
Some 33,812,183 objects have validated (non NULL) RV values in DR3. Among them some 995315 objects obey the input criteria. We recall them: RVs always between –1000 and +1000 km, brighter than 12, effective temperature between 3875 and 8125 K, and rv_normalised_gof above 4. Only the objects suspected of variability enter the orbital solution chain. Thus, an additional cut-off of rv_chisq_pvalue at 0.01 (kept below 0.01) further decreases the sample. A little more than 866 kilo-objects (866403) entered the present main NSS-SB1 processing. If we consider the SB1+SB2 processing, the total amount is 876644. The repartition of the results among the different types of solutions introduced in Sect. 3 is given in Table 1. The missing objects correspond to a lack of possible solutions. An important contribution comes from the objects having smaller than 10.
@@
After this preliminary main NSS-SB1 processing, the internal filtering is applied, as explained in Sect. 7.1. The corresponding numbers are given in Table 2 for the SB1-type solutions, in Table 3 for the SB1C-type solutions and finally, in Table 4 for the TrendSB1-type solutions. We give the effect of each criterion on the tallies. The internal filtering has been actually performed in the same sequence as given in the table. These tallies are however only indicative because they can be dependent on the order of the adopted filtering. Only the number of entering objects and the number of remaining objects are not sensitive to the adopted order. As explained in Sect. 2, some objects can enter the SB1 processing coming from the SB2 processing (see also the details in Damerdji et al. 2024). Table 5 deals with the very few objects coming from the SB2 chain. Due to the filtering on flags nr.11 and nr.12, these objects do not pervade the final SB-subcatalogue. No class is associated to these objects.
@@
We have now to estimate the number of solutions from our pipeline that were removed because they were combined and are now appearing in the EclipsingSpectro and AstroSpectroSB1 classes. Since we are only interested here on the disappearance of some solutions from our output due to the combination process, we can formally consider the tallies as if the combination were made after the internal filtering. The results of the combination mechanism for the 198885 SB1-type solutions are detailed in Table 5. The column Single corresponds to the solutions appearing in one class only whereas the column Duplicated refers to the objects appearing in two classes. All the 198885 objects are classified, except 2201 that will not be published due to a rejection during the process. Among these 2201 objects, 2107 correspond to a rejection at the level of the post-filtering which is detailed below. Some 94 other objects were rejected for other reasons. The majority of these 94 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue. The Orbital and Others classes are given for the sake of completeness but they are not further considered either in the present work nor, for some, in DR3. Only the first three classes of the Table 5 are of importance for the present work. We will in addition restrict the present analysis of the SB1-type solutions to the SB1 class. To illustrate the approach, we will further consider the case of EclipsingSpectro. According to Gaia Collaboration et al. (2023a), some 155 objects are in this class. Over these 155 objects, 1 was originally attributed a solution of type SB2C, one SB1C, 2 SB2 and 151 SB1. Among these 151 SB1 solutions coming out of our main processing, 79 were internally filtered out. Therefore, some 72 solutions were dispatched to EclipsingSpectro and did not reach the SB1 class.
@@
Finally, the TrendSB1-type solutions were compared with the astrometric channel results and 891 turned out to be redirected into the AstroSpectroSB1 class. They thus are not present in the TrendSB1 class, leading to an amount of 56808 objects. The details are explicited in Table 7. The majority of the 21 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue. In this table, we also counted the TrendSB1 objects that presented or not a significant period at the level of the main processing (flag nr.13 raised means no period found).
@@
After the combination process, a further step of filtering took place. This one was called post-filtering, and the motivation and practical details are described below in Sect. 10.1. The post-filtering was done on the basis of data coming from outside the RVS channel and is basically a process of blacklisting aiming at improving the purity of the catalogue. However, it should be made clear that it is partly constrained by the fact that the spectroscopic solutions are further used by the combiner, and thus we should be careful not to eliminate good solutions that could be useful downstream in the process. In the present section, we restrict ourselves to the effect of the applied post-filtering on the countings of the objects among the various classes. Table 6 details the effect on the number of objects considered in the important classes. The amount of objects before the post-filtering and after it are mentioned. The column after post-filtering gives the countings that are corresponding to the final DR3 catalogue.
@@ 8.2.1 Good results
In the present section, we introduce a few examples of the solution fitted to selected RV time-series. The selection is somewhat arbitrary and is essentially motivated by the illustration of the efficiency of the chain. All these illustrating good solutions are part of the DR3 spectroscopic orbit catalogue (SB-subcatalogue). Figure 54 illustrates various orbital solutions that are part of the SB1 class. A small subset of these SB1 solutions is also provided in Fig. 6. The displayed phase diagrams cover a palette of values for the period associated to the SB1 solution. It is a first positive point that the pipeline is able to treat such a large domain of periodicity. At mid to large period, we also present a few examples of highly eccentric orbits, up to . It is a matter of fact that a good fit to a very eccentric orbit necessitates a large number of data points in the time-series, particularly near the periastron. An homogeneous distribution in phase is another necessity.
@@
Since the vast majority of the analysed time-series covers a span of time around 800-1000 d, the solutions with short periods are extremely convincing since the large number of cycles folded in the phase diagram do not mitigate the goodness of fit and this can be associated to a high coherency of the periodicity. This conclusion holds when more than three cycles are covered (i.e. , or less). For larger periods, the coherency of the variations is not fully guaranteed by the quality of the goodness of fit. At larger values (), we are no more able to guarantee the coherency of the variation, since only one cycle is observed and this could correspond to a transitory event, not to a periodic variation. In any case, a tentative fit is still more indicative than no fit at all. However, the SB1-type solutions with such large up to (the adopted upper value allowed) should be considered as not secure as periodic solutions and the reported period is at the very best an indicative timescale. Indeed, for given periods between and 1.5 , the periods are not well determined and could even be potentially biased, due to the non-linear nature of the fit interacting with a possible bad phase coverage. This topic is further addressed in Sect. 9.4.1 and in Sect. 9.4.2. As illustrated below, the latter bias could induce biases on other parameters (see e.g. the case of the eccentricity in Fig. 64). The objects Gaia DR3 6422103351757549056, Gaia DR3 2509001872917567360, and Gaia DR3 5818903954144888320 are certainly concerned by this possible pitfall.
@@
It is interesting to notice that several objects in the sample correspond to a quasi-circular solution: the eccentricity is very small, even perhaps not significant. A perfect illustration (see Fig. 6) is the case of Gaia DR3 1955225454244655872, whose eccentricity is with a 1- error of 0.021. This object most probably corresponds to a circular orbit, but the adopted solution is not an SB1C but an SB1. This SB1-type solution is associated to a solution efficiency of 0.16 and is thus on the verge to be processed with an SB1C model. This is due to the fact that we adopted the philosophy of not enforcing a circular solution even if the eccentricity turned out not to be significantly different from zero. We only adopt a circular solution when the structure matrix is such that it could not be properly inverted in the classical way and thus a full SB1-type solution is not possible. We are consequently constrained to decrease the number of free parameters. In fact, Gaia DR3 1955225454244655872 benefited from a large number of observations (good transits) and this large number of data points allowed us to fit an SB1 solution even for a very small apparent value of the eccentricity. Another example of truly circular orbit could be e.g. Gaia DR3 515682154702230272. Starting from these examples, we encourage the user of the catalogue to understand that the classes SB1 and SB1C differ by the fitted model and NOT by the true nature of the object. For example, someone interested in objects in circular orbits, i.e. to study only genuine circular orbit should certainly consider both SB1 and SB1C classes. If a user wants to transform an SB1-type solutions into an SB1C one, they can certainly use the period, the centre-of-mass velocity and the semi-amplitude as given by the SB1-type solution. However, it should not be overlooked that the argument of periastron is thus modified, since the for SB1C solutions corresponds by definition to the time of maximum velocity. Therefore, the must be accordingly corrected for this change of . Otherwise, the SB1-type solution is applicable to objects in true circular orbits.
@@
Another interesting and enlightening case is the one of Gaia DR3 4983214300285468288. As can be seen from Fig. 6, the fit of the SB1-type solution is very good and no individual data point deviates significantly from the model. However, an expert eye (not the pipeline) will detect that around phase 0.65, the observed RVs are systematically grouped below the model and then above the modelled curve. Both groups of data points contain some 10 points and the offset could thus be significant (at least marginally) although the grouping is observed a posteriori. In any case, if this anomaly at phase 0.65 is real (and this will become clear in future releases), it is hardly explained in the framework of the Keplerian motion. The time corresponds to a phase where the primary is approaching the maximum velocity and thus the possible lines of the secondary are close to the maximum separation and should not influence the determination of the primary RVs (primary meaning here the star that dominates the combined spectrum). We are also (most probably) far from a possible eclipse, and this could not be due to a Rossiter-McLaughlin effect. The reported anomaly, if confirmed, will necessitate exotic effects to be explained in the context of a Keplerian model. It is worrying that several intrinsic variables (pulsating stars at least) could exhibit such a phenomenon at approximately this location in their RV curve. It is beyond the scope of the present paper to make a decision for this particular object but we would like to take this opportunity to highlight the following statement. It is not because a beautiful fit can be performed with a Keplerian model by the pipeline that the true nature of the object is the one of a binary. The membership to the SB1 or SB1C classes of solutions does not automatically imply the binarity of the object. Additional clues are necessary to converge to a sound classification. The binary nature of SB2 objects is much more convincing (although problems could also exist, see Damerdji et al. 2024). Concerning SB1 or SB1C objects, a conclusion could be drawn if, for example, the orbit is confirmed by the photometric or the astrometric channels (classes EclipsingSpectro or AstroSpectroSB1). The combined solutions are not part of the classes treated in the present paper. For the isolated spectroscopic orbital solutions reported here, we recommend at the very least, to build up a subcatalogue constituting some kind of gold sample, or to combine with additional data (e.g. positions in the HR diagram) in order to increase the veracity of the proposed classification.
@@
In Fig. 57, we exhibit a few examples of orbital solutions of the class SB1C. A still smaller subset is provided in Fig. 7. A general finding is the fact that most of the concerned time-series are characterised by a rather limited number of data points. They also correspond to objects characterised by . This is related to the decision we took to fit an SB1C-type solution only when the SB1 failed: we must be constrained to limit the amount of free parameters. We recall here that the convention has been adopted that, for SB1C-type solutions, the represents the time of passage at maximum velocity. We note from Fig. 57 that we have circular solutions for a few large period objects. Here also, the majority of the objects were not previously known to present a periodic behaviour in their RV.
The object Gaia DR3 3384500857079232512 has been observed with Kepler and a period d (or twice that value) has been detected (Armstrong et al. 2015). This does not seem to be coherent with our results (d). The solution of this discrepancy will certainly be automatically found in future releases thanks to the important increase of the number of data points in the time-series under analysis. This conclusion is supported by the simulations reported in Sect. 9.1.
@@
The solutions shown in Figs. 54, 57 and 58 are just illustrative and much more objects with solutions of similar quality are present in the SB-subcatalogue. This subcatalogue also contains solutions of slightly less good quality as well as solutions of intermediate quality. In the Appendix E, we provide a few examples of SB1-type solutions selected from the SB-subcatalogue at random but with around 0 (Fig. 59). This is to have fits of similar quality but for which the number of data points is lower (10 to 45) than in Fig. 6. Although the quality of these solutions is sometimes less evident, only very few could be easily questionable. Since we decided to push the software to its limit, solutions of more limited quality are equally present. For that reason, we encourage the general user to further filter the data. This is possible thanks to the quality factors that we deliver for each object (see Sect. 6.4). For a general statistical use, this should generate no problem, provided the user adopts the filtering to its particular application. We encourage the user to be very careful about the results on individual objects that could not benefit from the statistics. Some further information and advice can be found below in Sect. 11.2.
@@ 8.2.2 Problematic cases: multi-variability
In the framework of DR3, the SB1 pipeline is intended to search for orbital solutions. It is however designed to treat the cases which present only one type of variability (either SB-type with a single period or pure trends). If one object presents two types of variability, most of the time the algorithm could fail to find an adequate solution. We present here a typical case of badly handled data. Figure 9 shows the time-series of a star observed by Gaia RVS ( d). The data are unpublished and the solution is not part of the DR3 catalogue. The star is thus kept anonymous. The pipeline detected for this object a linear trend of about 1.15 km s-1 y-1. The solution corresponds to . This is easily understood because the RV curve exhibits extra-variations with a kind of oscillation with a period a little over 300 d. The variation is larger than the error bars and is thus significant. It necessarily induces a large value for coming from the fit of the trend, and the corresponding solution has thus been rejected from the DR3 catalogue. This object actually presents two superposed, different, types of variability. A trend that could be associated to a long period binary and, in addition, a roughly sinusoidal variation that could be due to an SB1 system in a triple system, or to an intrinsic RV variation or to something else. If, by chance, the pipeline would have first detected and reported an SB1 solution, the existing trend would have also implied a large associated to the SB1 fit and the solution would also have been rejected. This is clearly a weak point of the DR3 chain.

### v2

@@ 8 Results
DR3 is the first release providing orbital solutions for the RV time series. As such, the present results must be considered preliminary even if the performances are already good.
@@ 8.1 Tallies
Some 33 812 183 objects have validated (non NULL) RV values in DR3. Among them, some 995 315 objects obey the input criteria. We recall them: RVs always between –1000 and +1000 km, brighter than 12, effective temperature between 3875 and 8125 K, and rv_renormalised_gof above 4. Only the objects suspected of variability enter the orbital solution chain. Thus, an additional cut-off of rv_chisq_pvalue at 0.01 (kept below 0.01) further decreases the sample. A little more than 866 kilo-objects (866 403) entered the present main NSS-SB1 processing. If we consider the SB1+SB2 processing, the total amount is 876 644. The repartition of the results among the different types of solutions introduced in Sect. 3 is given in Table 1. The missing objects correspond to a lack of possible solutions. An important contribution comes from the objects having smaller than 10.
@@
After this preliminary main NSS-SB1 processing, the internal filtering is applied, as explained in Sect. 7.1. The corresponding numbers are given in Table 2 for the SB1-type solutions, in Table 3 for the SB1C-type solutions and finally, in Table 4 for the TrendSB1-type solutions. We give the effect of each criterion on the tallies. The internal filtering has been actually performed in the same sequence as given in the table. These tallies are however only indicative because they can be dependent on the order of the adopted filtering. Only the number of entering objects and the number of remaining objects are not sensitive to the adopted order. As explained in Sect. 2, some objects can enter the SB1 processing coming from the SB2 processing (see also the details in Damerdji et al. 2024). Table 5 deals with the very few objects coming from the SB2 chain. Due to the filtering on flags nr.11 and nr.12, these objects do not pervade the final SB subcatalogue. No class is associated to these objects.
@@
We have now to estimate the number of solutions from our pipeline that were removed because they were combined and are now appearing in the EclipsingSpectro and AstroSpectroSB1 classes. Since we are only interested here in the disappearance of some solutions from our output due to the combination process, we can formally consider the tallies as if the combination were made after the internal filtering. The results of the combination mechanism for the 198 885 SB1-type solutions are detailed in Table 5. The column marked ’Single’ corresponds to the solutions appearing in one class only, whereas the column marked ’Duplicated’ refers to the objects appearing in two classes. All the 198 885 objects are classified, except 2201 that will not be published due to a rejection during the process. Among these 2201 objects, 2107 correspond to a rejection at the level of the post-filtering which is detailed below. Some 94 other objects were rejected for other reasons. The majority of these 94 unpublished objects are corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB subcatalogue. The Orbital and Others classes are given for the sake of completeness but they are not further considered either in the present work nor, for some, in DR3. Only the first three classes of the Table 5 are of importance for the present work. We will in addition restrict the present analysis of the SB1-type solutions to the SB1 class. To illustrate the approach, we will further consider the case of EclipsingSpectro. According to Gaia Collaboration et al. (2023a), some 155 objects are in this class. Over these 155 objects, 1 was originally attributed a solution of type SB2C, 1 of type SB1C, 2 of type SB2, and 151 of type SB1. Among these 151 SB1 solutions coming out of our main processing, 79 were internally filtered out. Therefore, some 72 solutions were dispatched to EclipsingSpectro and did not reach the SB1 class.
@@
Finally, the TrendSB1-type solutions were compared with the astrometric channel results and 891 turned out to be redirected into the AstroSpectroSB1 class. They thus are not present in the TrendSB1 class, leading to an amount of 56 808 objects. The details are explicited in Table 7. The majority of the 21 unpublished objects are corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB subcatalogue. In this table, we also counted the TrendSB1 objects that presented or not a significant period at the level of the main processing (flag nr.13 raised means no period found).
@@
After the combination process, a further step of filtering took place. This one was called post-filtering, and the motivation and practical details are described below in Sect. 10.1. The post-filtering was done on the basis of data coming from outside the RVS channel and is basically a process of blacklisting aiming at improving the purity of the catalogue. However, it should be made clear that it is partly constrained by the fact that the spectroscopic solutions are further used by the combiner, and thus we should be careful not to eliminate good solutions that could be useful downstream in the process. In the present section we restrict ourselves to the effect of the applied post-filtering on the countings of the objects among the various classes. Table 6 details the effect on the number of objects considered in the important classes. The amount of objects before the post-filtering and after it are mentioned. The column after post-filtering gives the countings that are corresponding to the final DR3 catalogue.
@@ 8.2.1 Good results
Here, we introduce a few examples of the solution fitted to selected RV time series. The selection is somewhat arbitrary and is essentially motivated by the illustration of the efficiency of the chain. All these illustrating good solutions are part of the DR3 spectroscopic orbit catalogue (SB subcatalogue). Figure 54 illustrates various orbital solutions that are part of the SB1 class. A small subset of these SB1 solutions is also provided in Fig. 6. The displayed phase diagrams cover a palette of values for the period associated to the SB1 solution. It is a first positive point that the pipeline is able to treat such a large domain of periodicity. At mid to large period, we also present a few examples of highly eccentric orbits, up to . It is a matter of fact that a good fit to a very eccentric orbit necessitates a large number of data points in the time series, particularly near the periastron. An homogeneous distribution in phase is another necessity.
@@
Since the vast majority of the analysed time series covers a span of time around 800-1000 d, the solutions with short periods are extremely convincing since the large number of cycles folded in the phase diagram do not mitigate the goodness of fit and this can be associated to a high coherency of the periodicity. This conclusion holds when more than three cycles are covered (i.e. , or less). For larger periods, the coherency of the variations is not fully guaranteed by the quality of the goodness of fit. At larger values (), we are no more able to guarantee the coherency of the variation, since only one cycle is observed and this could correspond to a transitory event, not to a periodic variation. In any case, a tentative fit is still more indicative than no fit at all. However, the SB1-type solutions with such large up to (the adopted upper value allowed) should be considered as not secure as periodic solutions and the reported period is at the very best an indicative timescale. Indeed, for given periods between and 1.5 , the periods are not well determined and could even be potentially biased, due to the non-linear nature of the fit interacting with a possible bad phase coverage. This topic is further addressed in Sect. 9.4.1 and in Sect. 9.4.2. As illustrated below, the latter bias could induce biases on other parameters (see e.g. the case of the eccentricity in Fig. 64). The objects Gaia DR3 6422103351757549056, Gaia DR3 2509001872917567360, and Gaia DR3 5818903954144888320 are certainly concerned by this possible pitfall.
@@
It is interesting to notice that several objects in the sample correspond to a quasi-circular solution: the eccentricity is very small, even perhaps not significant. A perfect illustration (see Fig. 6) is the case of Gaia DR3 1955225454244655872, whose eccentricity is with a 1- error of 0.021. This object most probably corresponds to a circular orbit, but the adopted solution is not an SB1C but an SB1. This SB1-type solution is associated to a solution efficiency of 0.16 and is thus on the verge to be processed with an SB1C model. This is due to the fact that we adopted the philosophy of not enforcing a circular solution even if the eccentricity turned out not to be significantly different from zero. We only adopt a circular solution when the structure matrix is such that it could not be properly inverted in the classical way and thus a full SB1-type solution is not possible. We are consequently constrained to decrease the number of free parameters. In fact, Gaia DR3 1955225454244655872 benefited from a large number of observations (good transits) and this large number of data points allowed us to fit an SB1 solution even for a very small apparent value of the eccentricity. Another example of truly circular orbit could be e.g. Gaia DR3 515682154702230272. Starting from these examples, we encourage the user of the catalogue to understand that the classes SB1 and SB1C differ by the fitted model and NOT by the true nature of the object. For example, someone interested in objects in circular orbits, i.e. to study only genuine circular orbit, should certainly consider both SB1 and SB1C classes. If a user wants to transform an SB1-type solutions into an SB1C one, they can certainly use the period, the centre-of-mass velocity and the semi-amplitude as given by the SB1-type solution. However, it should not be overlooked that the argument of periastron is thus modified, since the for SB1C solutions corresponds by definition to the time of maximum velocity. Therefore, the must be accordingly corrected for this change of . Otherwise, the SB1-type solution is applicable to objects in true circular orbits.
@@
Another interesting and enlightening case is the one of Gaia DR3 4983214300285468288. As can be seen from Fig. 6, the fit of the SB1-type solution is very good and no individual data point deviates significantly from the model. However, an expert eye (not the pipeline) will detect that around phase 0.65, the observed RVs are systematically grouped below the model and then above the modelled curve. Both groups of data points contain some 10 points and the offset could thus be significant (at least marginally) although the grouping is observed a posteriori. In any case, if this anomaly at phase 0.65 is real (and this will become clear in future releases), it is hardly explained in the framework of the Keplerian motion. The time corresponds to a phase where the primary is approaching the maximum velocity and thus the possible lines of the secondary are close to the maximum separation and should not influence the determination of the primary RVs (primary meaning here the star that dominates the combined spectrum). We are also (most probably) far from a possible eclipse, and this could not be due to a Rossiter-McLaughlin effect. The reported anomaly, if confirmed, will necessitate exotic effects to be explained in the context of a Keplerian model. It is worrying that several intrinsic variables (pulsating stars at least) could exhibit such a phenomenon at approximately this location in their RV curve. It is beyond the scope of the present paper to make a decision for this particular object but we would like to take this opportunity to highlight the following statement. It is not because a beautiful fit can be performed with a Keplerian model by the pipeline that the true nature of the object is the one of a binary. The membership to the SB1 or SB1C classes of solutions does not automatically imply the binarity of the object. Additional clues are necessary to converge to a sound classification. The binary nature of SB2 objects is much more convincing (although problems could also exist; see Damerdji et al. 2024). Concerning SB1 or SB1C objects, a conclusion could be drawn if, for example, the orbit is confirmed by the photometric or the astrometric channels (classes EclipsingSpectro or AstroSpectroSB1). The combined solutions are not part of the classes treated in the present paper. For the isolated spectroscopic orbital solutions reported here, we recommend at the very least, to build up a subcatalogue constituting some kind of gold sample, or to combine with additional data (e.g. positions in the HR diagram) in order to increase the veracity of the proposed classification.
@@
In Fig. 57, we exhibit a few examples of orbital solutions of the class SB1C. A still smaller subset is provided in Fig. 7. A general finding is the fact that most of the concerned time series are characterised by a rather limited number of data points. They also correspond to objects characterised by . This is related to the decision we took to fit an SB1C-type solution only when the SB1 failed: we must be constrained to limit the amount of free parameters. We recall here that the convention has been adopted that, for SB1C-type solutions, the represents the time of passage at maximum velocity. We note from Fig. 57 that we have circular solutions for a few large period objects. Here also, the majority of the objects were not previously known to present a periodic behaviour in their RV.
The object Gaia DR3 3384500857079232512 has been observed with Kepler and a period d (or twice that value) has been detected (Armstrong et al. 2015). This does not seem to be coherent with our results (d). The solution of this discrepancy will certainly be automatically found in future releases thanks to the important increase of the number of data points in the time series under analysis. This conclusion is supported by the simulations reported in Sect. 9.1.
@@
The solutions shown in Figs. 54, 57, and 58 are just illustrative and much more objects with solutions of similar quality are present in the SB subcatalogue. This subcatalogue also contains solutions of slightly less good quality as well as solutions of intermediate quality. In Appendix E, we provide a few examples of SB1-type solutions selected from the SB subcatalogue at random but with around 0 (Fig. 59). This is to have fits of similar quality but for which the number of data points is lower (10 to 45) than in Fig. 6. Although the quality of these solutions is sometimes less evident, only very few could be easily questionable. Since we decided to push the software to its limit, solutions of more limited quality are equally present. For that reason, we encourage the general user to further filter the data. This is possible thanks to the quality factors that we deliver for each object (see Sect. 6.4). For a general statistical use, this should generate no problem, provided the user adopts the filtering to its particular application. We encourage the user to be very careful about the results on individual objects that could not benefit from the statistics. Some further information and advice can be found below in Sect. 11.2.
@@ 8.2.2 Problematic cases: multi-variability
In the framework of DR3, the SB1 pipeline is intended to search for orbital solutions. It is however designed to treat the cases which present only one type of variability (either SB-type with a single period or pure trends). If one object presents two types of variability, most of the time the algorithm could fail to find an adequate solution. We present here a typical case of badly handled data. Figure 9 shows the time series of a star observed by Gaia RVS ( d). The data are unpublished and the solution is not part of the DR3 catalogue. The star is thus kept anonymous. The pipeline detected for this object a linear trend of about 1.15 km s-1 y-1. The solution corresponds to . This is easily understood because the RV curve exhibits extra-variations with a kind of oscillation with a period a little over 300 d. The variation is larger than the error bars and is thus significant. It necessarily induces a large value for coming from the fit of the trend, and the corresponding solution has thus been rejected from the DR3 catalogue. This object actually presents two superposed, different, types of variability. A trend that could be associated to a long period binary and, in addition, a roughly sinusoidal variation that could be due to an SB1 system in a triple system, or to an intrinsic RV variation or to something else. If, by chance, the pipeline would have first detected and reported an SB1 solution, the existing trend would have also implied a large associated to the SB1 fit and the solution would also have been rejected. This is clearly a weak point of the DR3 chain.